OpenStudy (anonymous):

Hello, who know the midpoint of D(1, 2) and E(–3, 6)

OpenStudy (kobeni-chan):

You find the mean of each variable. 1 + (-3) = -2 -2/2 = -1 2 + 6 =8 8/2 = 4 The midpoint is (-1,4)

OpenStudy (anonymous):

thank you!!!! do you know this one? D(–5, 2) and E(–13, 0)

OpenStudy (anonymous):

(5,2) (13,0) what's the mindpoint?